OpenStudy (anonymous):

PLEASE HELP : If f(x) = 5x – 4 and g(x) = 6x + 3, what is f(g(x)) ?

OpenStudy (anonymous):

substitute g(x) for x in f(x) \[f(g(x))=5(6x+3)-4\] simplify, why do you disappear immediately after you post a question

OpenStudy (anonymous):

i dont disappear ?

OpenStudy (anonymous):

and im still a bit lost

OpenStudy (anonymous):

\[f(g(x))=f(6x+3) = 5(6x+3)-4\]

OpenStudy (anonymous):

they want you to use the output of g(x) as the input for f(x)

OpenStudy (anonymous):

start with \[f(g(x))\] replace \[g(x)\] by \[6x+3\] then where youi see and 'x" in \[f(x)\] replace it by \[ 6x+3\]

OpenStudy (anonymous):

f(x) = 5x - 4 So, look at that x Now, g(x) = 6x + 3 f( g(x) ) g(x) function goes in place of "x" in f(x) = 5x-4 Thus, f( g(x) ) = 5 (6x+3) - 4 And as you solve, f( g(x) ) = 60x + 15 - 4 = 60x + 11

OpenStudy (anonymous):

yeah but my answer choices are 30x – 21 30x – 1 30x + 11 30x + 12

OpenStudy (anonymous):

lol my bad i multiplied it wrong. it's 30x + 11

OpenStudy (anonymous):

okay thank you (: